# AdaMesh: Personalized Facial Expressions and Head Poses for Adaptive Speech-Driven 3D Facial Animation

### v2

@@ 1 Introduction
Speech-driven 3D facial animation has been widely explored and attracted increasing interest from both academics and industry. This technology has great potential in virtual reality, film production, and game creation. Most previous works focus on improving synchronization between speech and lip movements [5, 23, 8, 34], however, they have neglected the personalized talking style, including the facial expressions and head poses. The facial animation with expressive facial expressions and diverse head poses contributes to a more vivid virtual character and grants users with a better interaction experience.
Figure 1: The overview of AdaMesh. The expression adapter and pose adapter generate personalized facial expressions and head poses with the given speech signal and reference talking styles.
@@
Recently, some works [35, 28, 26] attempt to model the person-specific talking style by fine-tuning or adapting modules. In the real application, only a limited amount of video clips, even shorter than 1 minute, are provided by target users to capture the personalized talking style. These methods thus meet several challenges. 1) The few adaptation data probably induces catastrophic forgetting [37] for the pre-trained model and easily leads to the overfitting problem [39]. Specifically for the facial expressions, after the adaptation, the lip synchronization and richness of expressions significantly decrease for the unseen speech inputs. 2) Speech is a weak control signal for the head poses [12, 38]. Adaption with few data or learning mappings on such weak signals leads to averaged generation. The predicted head poses are in lack of diversity and have a smaller variation of movements. 3) The facial expressions and head poses are modeled by a one-size-fits-all approach [38, 35], neglecting the intrinsic characteristics of the data.
In this work, to address these challenges, we propose AdaMesh as illustrated in Fig. 1, an adaptive speech-driven 3D facial animation approach. In consideration of the distinct data characteristics of facial expressions and head poses, we devise two specialized adapters—an expression adapter and a pose adapter—tailored to effectively capture the person-specific talking styles, eschewing the one-size-fits-all approach.
@@
To tackle the catastrophic forgetting and overfitting problem, we devise an expression adapter built upon the low-rank adaptation technique, LoRA[14], for its high data efficiency in the few-shot model adaptation. To model the multi-scale temporal dynamics in facial expressions, we extend the vanilla LoRA to the convolution operation and further propose a novel mixture-of-LoRA (MoLoRA) strategy, featured by employing LoRA with different rank sizes.
@@
As for the averaged pose generation problem, we recognize it as a repercussion of generating poses via motion regression [34] and a mismatch of the adapted pose style. Hence we formulate the pose generation task as a generative one, for which we employ a vector quantized-variational autoencoder (VQ-VAE) [30] and a generative pre-trained Transformer (PoseGPT) network [32] as the backbone of the pose adapter. The head poses are mostly associated with the conveyance of negation, affirmation, and turnaround in semantics. To balance the strong and weak semantic associations in head pose patterns, we propose the retrieval strategy for the pose style adaptation with a novel semantic-aware pose style matrix.
The main contributions can be summarized as follows:
@@
  • We propose AdaMesh, which learns personalized facial expressions and head poses for speech-driven 3D facial animation with limited adaptation data.
@@
  • We are the first to introduce LoRA to the 3D facial animation task and propose MoLoRA to efficiently capture the multi-scale temporal dynamics in facial expressions.
@@
  • We propose a pose adapter, featured by a semantic-aware pose style matrix and a simple retrieval strategy for diverse pose generation without fine-tuning any parameters.
@@ 2 Related Work
Speech-Driven 3D Facial Animation. The recent speech-driven 3D facial animation methods can be divided into two categories according to the format of animation. The first category of methods [5, 23, 8, 34, 28] directly predicts vertices of the face from speech and renders the vertices to meshes for visualization. The vertice data is collected through professional 3D scans of human heads [9, 5, 16], which is expensive and time-consuming. The methods trained on such data usually have authentic lip movements. Another category of methods predicts parameters of expressions [6] or blendshapes [1, 21]. These parameters can be reconstructed from the scanned vertices or 2D face videos through a parametric face model [16, 21]. However, the decreasing quality caused by the reconstruction usually leads to lower speech-lip synchronization. The latest methods that consider personalized style mostly adopt expression parameters for style modeling, since the scanned vertices are in the neutral style and lack rich expressions. Emotalk [21] and EMOTE [6] are related works, which disentangle content and emotion to generate emotional talking styles. But they adopt private or unavailable reconstruction pipelines. The most related work is Imitator [28], which generates meshes and leverages the expressions from 2D video for style adaptation.
@@
Head Poses Generation. To the best of our knowledge, existing speech-driven 3D facial animation methods neglect the head poses, since the scanned vertices have no head movements. Some portrait-realistic talking face generation methods attempt to model the head poses. A typical way is to copy the head pose sequence from a real video recording [29, 35, 2], which can definitely generate realistic head poses. However, these methods disregard the correlation between speech and head pose, and the repetitive and non-engaging head poses create unnaturalness during long periods of interaction. The rapidly growing research topic, gesture generation from speech [41], also justifies the importance of predicting motions from speech rather than utilizing a template motion sequence. Another way to predict head poses from speech also attracts many works [19, 38, 40]. However, these methods can only generate head poses with specific styles, and cannot realize flexible style adaptation. Our approach could be regarded as bridging the advantages of the aforementioned two ways. The quantized pose space learned by VQ-VAE ensures realism, while PoseGPT predicts natural but not repetitive poses.
@@
Adaptation Strategy. Efficient adaptation strategies of pre-trained models, including adapter-based [22], prefix-based [17], and LoRA [14] techniques, have been applied to various downstream applications. These strategies achieve advanced performance without fine-tuning all the model parameters. In the field of image synthesis [24], LoRA serves as a plug-and-play extension for the pre-trained model to bring tremendous and specific styles to the images. This work is inspired by this idea and intends to bring the person-specific style to the pre-trained networks that predict facial expressions from speech. Note that we refer to the models proposed in Sec. 3.1 and Sec. 3.2 as the adapters. It is different from the concept of adapter-based strategy [22] mentioned above.
@@
Remark. Our work differs from previous works by the enhanced data efficiency in style adaptation and overcomes the challenges of catastrophic forgetting for facial expressions and averaged pose generation.
Figure 2: The overview of expression adapter. (a) The patches on each encoder and decoder denote the MoLoRA parameters which are added to the pre-trained modules after adaptation. (b) A brief illustration of Conformer [15] (c) demonstrates that MoLoRA is applied to the convolution operator. It is also applied to the linear layers in the proposed expression adapter.
@@ 3 AdaMesh
In this section, we propose AdaMesh, a data-efficient speech-driven 3D facial animation approach tailored for generating personalized facial expressions and head poses. Given the facial expression sequence and head pose sequence extracted from a reference video, together with a driving speech, the expression adapter (Sec. 3.1) is adapted with the specific expression style and outputs rich facial expressions; while the pose adapter (Sec. 3.2) derives a semantic-pose aware style matrix, and outputs diverse head poses with VQ-VAE and PoseGPT. The generated facial expressions and head poses are combined by a parametric face model [16] and rendered to mesh images.
@@ 3.1 Expression Adapter
Audio Encoder & Identity Encoder. To obtain robust semantic information from speech, the self-supervised speech representation model HuBERT  [13] is leveraged to extract features from raw speech. These speech features are then sent into three Conformer layers [15]. Previous methods usually represent the identity with the one-hot label [34, 28], leading to the bad generalization on unseen speaker identity. We utilize the identity parameter extracted by the parametric face model [16, 10]. The identity sequence in a data sample is averaged along the time dimension to represent the speaker identity and then sent into three convolutional layers with conditional layer normalization [3].
@@
Style Encoder and Expression Decoder. To provide the talking style that cannot be directly predicted from the semantic speech information, the expression sequence is utilized as a residual condition to the expression prediction. The sequence is compressed to a compact style embedding after two Conformer layers and then concatenated with the encoded speech and identity features into the decoder. The decoder also consists of three Conformer layers, which are stacked in a residual manner [36] rather than cascadation to realize coarse-to-fine prediction.
Figure 3: The overview of pose adapter. (a) Training of VQ-VAE. (b) PoseGPT. (c) The derivation of the semantic-aware pose style matrix and the retrieval strategy. In the training of the PoseGPT, the pose style embedding is the assigned one-hot label for each sample. $S$ denotes the semantic-aware pose style matrix. More details about the training and inference can be found in the supplementary materials.
@@
MoLoRA. We focus on the low-rank training techniques and propose MoLoRA to equip the expression adapter with the inductive bias of new data. As it is known, the vanilla LoRA [14] injects trainable rank decomposition matrices into each linear layer of the Transformer architecture. Since the convolutional layer can be viewed as multiple linear transformations with sliding receptive windows, we extend LoRA to the convolution operator for more flexible trainable parameters. The facial expressions contain multi-scale temporal dynamics, where the coarse-grained global features (e.g., emotion) and fine-grained facial details (e.g., speech-sync lip movements, muscle wrinkles) inherent in the expression sequence. We empirically found that the vanilla LoRA with a small rank size tends to preserve global features and a bigger rank size captures more fine-grained dynamics. Hence we propose MoLoRA to mix $N$ LoRA with different rank sizes $r_{i}$. Supposing the pretrained weight matrix of 1D convolution $W_{0}\in\mathbb{R}^{m\times n\times k}$, we define the low-rank product $\Delta W^{i}=W_{B}^{i}W_{A}^{i}$, where $W_{B}^{i}\in\mathbb{R}^{m/r_{i}\times k,r_{i}\times k}$ and $W_{A}^{i}\in\mathbb{R}^{r_{i}\times k,n\times r_{i}}$. The weights $W$ after adaptation merge the pre-trained and MoLoRA weights, which can be defined as:
$\vspace{-1pt}W=W_{0}+\Delta W=W_{0}+\sum_{i=0}^{N}W_{B}^{i}W_{A}^{i}.\vspace{-1pt}$ (1)
@@
There is a concurrent work [18] that investigates the efficiency of LoRA. However, it stacks multiple LoRA of a single rank size, while MoLoRA focuses on modeling the multi-scale characteristic in facial expressions.
@@
Procedure of the Expression Adapter. The pipeline of the expression adapter is listed in Alg. 1. Note that the inference stage has a slight difference with the training. The style embedding is obtained from another expression sequence rather than the ground truth. However, this mismatch has no impact on the generation, since the style embedding learns a global talking style for the expression sequence rather than the semantic speech information.
Algorithm 1 Pipeline of Expression Adapter.
@@
  Pretrain: Train the Expression Adapter $\theta_{E}$ with source training data. The expression sequence for the style encoder is the same with the prediction of the expression adapter.
@@
  Adaptation: Update all low-rank matrices $\theta_{MoLoRA}$ for convolution and linear layers in the expression adapter with the adaptation data.
@@
  Inference: Merge $\theta_{E}$ and $\theta_{MoLoRA}$. Given a driving speech, and an expression sequence from the adaptation data as the input for the style encoder, the expression adapter generates expressions that synchronize with the speech and reenact the facial style of the adaptation data.
@@ 3.2 Pose Adapter
To overcome the averaged generation problem in pose style adaptation, we propose the pose adapter and formulate the adaptation as a simple but efficient retrieval task instead of fine-tuning modules. As illustrated in Fig. 3, we first learn a discrete latent codebook $\mathcal{Z}=\{\boldsymbol{z_{k}}\in\mathbb{R}^{d_{\boldsymbol{z}}}\}_{k=1}^{M}$ for head poses with the novel sequence-encoding VQ-VAE [25], and then predicts the discrete head pose codes from the driving speech with PoseGPT. The VQ-VAE training requires no speech input, so various video datasets of a larger amount can be utilized. This design enables the coverage of a wide range of pose styles, which is well-suited for learning a realistic pose manifold  [20]. A semantic-aware pose style matrix is derived for each pose sample in the training sample to create a database. We retrieve the nearest matrix in this database as the adapted style matrix in the adaptation stage. The simple retrieval strategy for adaptation is based on the assumption that the possible pose styles have been covered in the VQ-VAE learning with the large-scale dataset.
@@
Discrete Pose Space. Given head pose sequence in continuous domain $\boldsymbol{\Psi}=[\boldsymbol{\psi}_{1},\ldots,\boldsymbol{\psi}_{T}]\in\mathbb{R}^{T\times 3}$, the VQ-VAE encoder, which consists of 1D convolutional layers, firstly encodes it into context-aware features $\hat{\boldsymbol{Z}}\in\mathbb{R}^{T^{\prime}\times d_{\boldsymbol{z}}}$, where $T^{\prime}=T/w$ denotes the frame numbers of the downsampled features. Then, we obtain the quantized sequence $\boldsymbol{Z_{q}}$ via an element-wise quantization function $Q$ that maps each element in $\boldsymbol{Z_{q}}$ to its closest codebook entry:
$\vspace{-1pt}\boldsymbol{Z_{q}}=Q(\hat{\boldsymbol{Z}}):=(\underset{\boldsymbol{z_{k}}\in\mathcal{Z}}{\arg\min}\|\hat{\boldsymbol{z}}_{t}-\boldsymbol{z}_{k}\|_{2})\in\mathbb{R}^{T^{\prime}\times d_{\boldsymbol{z}}}.\vspace{-1pt}$ (2)
@@
$\boldsymbol{Z_{q}}$ is then reconstructed to continuous poses by VQ-VAE decoder, which consists of three 1D convolutional layers.
@@
PoseGPT. With the learned discrete pose space, we build a PoseGPT network [32, 25] that maps the driving speech and pose style embedding into the discrete codes of the target poses. The HuBERT feature and mel spectrogram of the driving speech are utilized for speech representation to consider both speech semantics and speech prosody (e.g., stress, tone) [1] information. We assign each pose sample in the training dataset with a one-hot label as the pose style embedding. The speech representation and pose style embedding are sent into PoseGPT, which consists of six cross-conditional Transformer layers [32]. PoseGPT predicts discrete codes autoregressive by selecting the output with the highest probability at each step. The predicted codes are finally sent into the VQ-VAE decoder to obtain the continuous head pose sequence. We train PoseGPT using the teacher-forcing training scheme with the cross-entropy loss.
@@
Semantic-Aware Pose Style Matrix. Given a speech utterance of $T$ frames and the corresponding head pose sequence, HuBERT [13] assigns each speech frame into one of 512 clusters, while the VQ-VAE encoder transforms the pose sequence into the latent sequence $\hat{\boldsymbol{Z}}$. Thus, each frame in $\hat{\boldsymbol{Z}}$ is labeled with a cluster which contains the speech semantic information. We calculate the semantic-aware pose style matrix $S\in\mathbb{R}^{512\times d_{z}}$ for each pose sequence as:
$\vspace{-1pt}S_{j}=\sum_{i=1}^{T^{\prime}}\hat{\boldsymbol{Z}}_{i}\cdot\delta(L_{i},j)/\sum_{i=1}^{T^{\prime}}\delta(L_{i},j),j=[1,\dots,512],\vspace{-1pt}$ (3)
@@
where $L_{i}$ denotes the cluster label for $i$ frame of $\hat{\boldsymbol{Z}}$ and $\delta$ is the Kronecker symbol that outputs 1 when the inputs are equal else 0. This equation aggregates all the poses which have the same semantic unit of HuBERT. Since the clustering step in HuBERT [13] captures the semantic information in acoustic units and the VQ-VAE encoder models the temporal pose style information, the derived $S$ represents the pose style that considers semantics.
@@
Procedure of the Pose Adapter. The pipeline of the pose adapter is listed in Alg. 2. Note that the adaptation stage involves no parameter updating.
Algorithm 2 Pipeline of Pose Adapter.
@@
  Pretrain: For each pose sequence in the source training dataset, assign it with a one-hot pose style embedding, and calculate its semantic-aware pose style matrix with its corresponding speech. Establish the database for all pose style matrices and their paired pose style embedding. Train the VQ-VAE and PoseGPT with the source training dataset.
@@
  Adaptation: Given an unseen pair of reference speech and pose sequence, calculate the semantic-aware pose style matrix for the reference data. Find the nearest pose style matrix in the established database by computing the L1 distance, and adopt the corresponding pose style embedding for inference.
@@
  Inference: Given a driving speech, and utilizing the retrieved pose style embedding, the pose adapter generates continuous head poses with the pre-trained PoseGPT and the VQ-VAE decoder.

### v3

@@ 1 Introduction
Speech-driven 3D facial animation has been widely explored and attracted increasing interest from both academics and industry. This technology has great potential in virtual reality, film production, and game creation. Most previous works focus on improving synchronization between speech and lip movements (Cudeiro et al., 2019; Richard et al., 2021; Fan et al., 2022; Xing et al., 2023), however, they have neglected the personalized talking style, including the facial expressions and head poses. The facial animation with expressive facial expressions and diverse head poses contributes to a more vivid virtual character and grants users with a better interaction experience.
@@
Recently, some works (Ye et al., 2023; Thambiraja et al., 2023) attempt to model the person-specific talking style by fine-tuning or adapting modules. In the real application, only a limited amount of video clips, even shorter than 1 minute, are provided by target users to capture the personalized talking style. These methods thus meet several challenges. 1) The few adaptation data probably induces catastrophic forgetting (Zhai et al., 2023) for the pre-trained model and easily leads to the overfitting problem (Zhang et al., 2022). Specifically for the facial expressions, after the adaptation, the lip synchronization and richness of expressions significantly decrease for the unseen speech inputs. 2) Speech is a weak control signal for the head poses (Henter et al., 2020; Zhang et al., 2021). Adaption with few data or learning mappings on such weak signals leads to averaged generation. The predicted head poses are in lack of diversity and have a smaller variation of movements.
@@
In this work, to address these two challenges, we propose an adaptive speech-driven 3D facial animation approach, AdaMesh, consisting of an expression adapter and a pose adapter. To tackle the catastrophic forgetting and overfitting problem, we devise an expression adapter built upon the low-rank adaptation technique, LoRA(Hu et al., 2022), for its high data efficiency in few-shot model adaptation. Since the facial expressions contain temporal dynamics, we propose a novel mixture-of-LoRA (MoLoRA) strategy, featured by employing LoRA with different rank sizes to model the multi-scale dynamic facial details.
@@
As for the averaged pose generation problem, we recognize it as a repercussion of generating poses via motion regression (Xing et al., 2023) and a mismatch of the adapted pose style. Hence we formulate the pose generation task as a generative one, for which we employ a vector quantized-variational autoencoder (VQ-VAE) (van den Oord et al., 2018) and a generative pre-trained Transformer (Pose GPT) network (Vaswani et al., 2017) for the backbone of the pose adapter. The head poses are mostly associated with the conveyance of negation, affirmation, and turnaround in semantics. To consider and balance both strong and weak semantic associations in head pose patterns, we propose the retrieval strategy for the pose style adaptation with a novel semantic-aware pose style matrix.
The main contributions of our work can be summarized as follows:
@@
  • We propose AdaMesh, which is the first work that takes both facial expressions and head poses into account in personalized speech-driven 3D facial animation.
@@
  • We are the first to introduce LoRA to the 3D facial animation task and propose MoLoRA to efficiently capture the multi-scale temporal dynamics in facial expressions from limited data.
@@
  • We propose a pose adapter, featured by a semantic-aware pose style matrix and a simple retrieval strategy for more diverse pose generation without fine-tuning any parameters.
@@ 2 Related Work
Speech-Driven 3D facial Animation. The recent speech-driven 3D facial animation methods can be divided into two categories according to the format of animation. The first category of methods (Cudeiro et al., 2019; Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Thambiraja et al., 2023) directly predicts vertices of the face from speech and renders the vertices to meshes for visualization. The vertice data is collected through professional 3D scans of human heads (Fanelli et al., 2010; Cudeiro et al., 2019; Li et al., 2017), which is expensive and time-consuming. The scanned data precisely describes the facial movements. The methods trained on such data usually have authentic lip movements. Another category of methods predicts parameters of expressions (Daněček et al., 2023) or blendshapes (Chen et al., 2022; Peng et al., 2023). These parameters can be reconstructed from the scanned vertices or 2D face videos through a parametric face model (Li et al., 2017; Peng et al., 2023). However, the decreasing quality caused by the reconstruction usually leads to lower speech-lip synchronization. The latest methods that take personalized style into consideration mostly adopt expression parameters for style modeling, since the scanned vertices are in the neutral style and lack rich expressions. Emotalk (Peng et al., 2023) and EMOTE (Daněček et al., 2023) are related works, which disentangle content and emotion to generate emotional talking style. But they adopt private or unavailable reconstruction pipelines. The most related work is Imitator (Thambiraja et al., 2023), which generates meshes and leverages the expressions from 2D video for style adaptation.
@@
Head Poses Generation. To the best of our knowledge, existing speech-driven 3D facial animation methods neglect the head poses, since the scanned vertices have no head movements. Some portrait-realistic talking face generation methods attempt to model the head poses. A typical way is to copy the head pose sequence from a real video recording (Thies et al., 2020; Ye et al., 2023; Chen et al., 2023), which can definitely generate realistic head poses. However, these methods disregard the correlation between speech and head pose, and the repetitive and non-engaging head poses create unnaturalness during long periods of interaction. The rapidly growing research topic, gesture generation from speech (Zhu et al., 2023), can also justify the importance of predicting head poses from speech rather than utilizing a template pose sequence. Another way is to predict head poses from speech (Lu et al., 2021; Zhang et al., 2021). The most effective method for head poses is FACIAL (Zhang et al., 2021), which proposes an implicit attribute learning framework to generate natural head poses. However, FACIAL sometimes generates static head poses and lacks pose diversity. Moreover, these methods can only generate head poses for a specific person, and cannot realize flexible style adaptation. Our approach could be regarded as bridging the advantages of the aforementioned two ways. The quantized pose space learned by VQ-VAE ensures realism, while the prediction step outputs natural but not repetitive pose sequence.
@@
Adaptation Strategy. Efficient adaptation strategies of pre-trained models, including adapter-based (Rebuffi et al., 2017), prefix-based (Li & Liang, 2021), and LoRA (Hu et al., 2022) techniques, have been applied to various downstream applications. These strategies achieve advanced performance without fine-tuning all the model parameters. In the field of image synthesis (Rombach et al., 2022), LoRA serves as a plug-and-play extension for the pre-trained model to bring tremendous and specific styles to the images. This work is inspired by this idea and intends to bring the person-specific style to the pre-trained networks that predict facial expressions from speech. A similar idea in speech synthesis (Hsieh et al., 2022) manages to clone the voice of a new speaker with LoRA. Note that we refer to the models proposed in Sec. 3.1 and Sec. 3.2 as the adapters. It is different from the concept of adapter-based strategy Rebuffi et al. (2017) mentioned above.
@@
Remark. Our work differs from previous works by the enhanced data efficiency in style adaptation and overcomes the challenges of catastrophic forgetting for facial expressions and averaged pose generation. To the best of our knowledge, we are the first to simultaneously take facial expressions and head poses into account in speech-driven 3D facial animation.
@@ 3 AdaMesh
In this section, we propose AdaMesh, a data-efficient speech-driven 3D facial animation approach tailored for generating personalized facial expressions and head poses. Given the extracted facial expression sequence and head pose sequence from a reference video, together with a driving speech, the expression adapter (Sec. 3.1) is adapted with a specific expression style and outputs rich facial expressions, while the pose adapter (Sec. 3.2) derives a semantic-pose aware style matrix and outputs diverse head poses. The generated facial expressions and head poses are combined by a parametric face model (Li et al., 2017) and rendered to mesh images.
Figure 1: The overview of expression adapter. (a) The patches on each encoder and decoder denote the MoLoRA parameters which are added to the pre-trained modules after adaptation. (b) A brief illustration of Conformer (Li et al., 2021) (c) demonstrates that MoLoRA is applied to the convolution operator. It is also applied to the linear layers in the proposed expression adapter.
@@ 3.1 Expression Adapter
Audio encoder. To obtain robust semantic information from speech, we leverage HuBERT (Hsu et al., 2021), a self-supervised speech representation model, to extract features from raw speech. Unlike Wav2Vec (Baevski et al., 2020) used in previous methods (Fan et al., 2022; Xing et al., 2023), HuBERT can preserve more semantic information and provide better cross-lingual ability. The speech features are downsampled to the same length with expression sequence using convolutional layers with stride 2 and then sent into three Conformer layers (Li et al., 2021).
@@
Identity encoder. Previous methods usually represent the identity with the one-hot label (Xing et al., 2023; Thambiraja et al., 2023) in the animation prediction. They show bad generalization on unseen speaker identity. We utilize the identity parameter extracted by the parametric face model (Li et al., 2017; Feng et al., 2021). The identity sequence in a data sample is averaged along the time dimension to represent the speaker identity and then sent into three convolutional layers with conditional layer normalization (Chen et al., 2021).
@@
Style encoder and Expression Decoder. To provide the talking style which cannot be directly predicted from the semantic speech information, the expression sequence is utilized as a residual condition to the expression prediction. The sequence is compressed to a compact style embedding after two Conformer layers and then concatenated with the encoded speech and identity features into the decoder. The decoder also consists of three Conformer layers, which are stacked in a residual manner rather than cascadation. It is inspired by the residual vector quantization (Zeghidour et al., 2021), which can realize coarse-to-fine prediction.
@@
MoLoRA. We focus on the low-rank training techniques and propose MoLoRA to equip the expression adapter with the inductive bias of new data. As it is known, the vanilla LoRA (Hu et al., 2022) injects trainable rank decomposition matrices into each linear layer of the Transformer architecture. Since the convolutional layer can be viewed as multiple linear transformations with sliding receptive windows, we extend LoRA to the convolution operator for more flexible trainable parameters. The facial expressions contain multi-scale temporal dynamics, where the coarse-grained expression style (e.g., emotion) and fine-grained facial details (e.g., speech-sync lip movements, muscle random wrinkles) inherent in the expression sequence. We empirically found that the vanilla LoRA with only a single rank size filters out the multi-scale dynamics and tends to produce over-smoothed facial expressions. Hence we propose MoLoRA to mix $N$ LoRA with different rank sizes $r_{i}$. Supposing the pretrained weight matrix of 1D convolution $W_{0}\in\mathbb{R}^{m\times n\times k}$, we define the low-rank product $\Delta W^{i}=W_{B}^{i}W_{A}^{i}$, where $W_{B}^{i}\in\mathbb{R}^{m/r_{i}\times k,r_{i}\times k}$ and $W_{A}^{i}\in\mathbb{R}^{r_{i}\times k,n\times r_{i}}$. The weights $W$ after adaptation merge the pre-trained and MoLoRA weights, which can be defined as:
$W=W_{0}+\Delta W=W_{0}+\sum_{i=0}^{N}W_{B}^{i}W_{A}^{i}.$ (1)
@@
There is a concurrent work (Lialin et al., 2023) that investigates the efficiency in LoRA. However, it stacks multiple LoRA of a single rank size, while MoLoRA focuses on modeling the multi-scale characteristic in facial expressions.
@@
Procedure of the expression adapter. The pretrain, adaptation, and inference procedures of the expression adapter are listed in Alg. 1. Note that the inference stage has a slight difference with the training. The style embedding is obtained from another expression sequence rather than the ground truth. However, this mismatch has no impact on the generation, since the style embedding learns a global talking style for the expression sequence rather than the semantic speech information.
Algorithm 1 Pretrain, Adaptation, and Inference of Expression Adapter.
@@
1:  Pretrain: Train the Expression Adapter $\theta_{E}$ with source training data. The expression sequence for the style encoder is the same with the prediction of the expression adapter.
@@
2:  Adaptation: Update all low-rank matrices $\theta_{MoLoRA}$ for convolution and linear layers in the expression adapter with the adaptation data.
@@
3:  Inference: Merge $\theta_{E}$ and $\theta_{MoLoRA}$. Given a driving speech, and using an expression sequence from the adaptation data as the input for the style encoder, the expression adapter generates expressions that synchronize with the speech and reenact the facial style of the adaptation data.
Figure 2: The overview of pose adapter. (a) Training of VQ-VAE. (b) Pose GPT. (c) The derivation of the semantic-aware pose style matrix and the retrieval strtegy. In the training of the Pose GPT, the pose style embedding is the assigned one-hot label. $S$ denotes the semantic-aware pose style matrix.
@@ 3.2 Pose Adapter
To overcome the averaged generation problem in pose style adaptation, we propose the pose adapter and formulate the adaptation as a simple but efficient retrieval task instead of fine-tuning modules. As illustrated in Fig. 2, we first learn a discrete latent codebook $\mathcal{Z}=\{\bm{z_{k}}\in\mathbb{R}^{d_{\bm{z}}}\}_{k=1}^{M}$ of head poses with the novel sequence-encoding VQ-VAE (Siyao et al., 2022), and predicts the speech-relevant head poses with Pose GPT. The VQ-VAE training requires no speech input, so various video datasets of a larger amount can be utilized. This design enables the coverage of a wide range of poses, which is well-suited for learning a realistic pose manifold (Ng et al., 2022). A semantic-aware pose style matrix is derived for each pose sample in the training sample to create a database. We retrieve the nearest matrix in the pose style matrix database as the adapted style matrix in the adaptation stage. The simple retrieval strategy for adaptation is based on the assumption that the possible pose styles are covered in the VQ-VAE learning with the large-scale dataset.
@@
Discrete Pose Space. Given a head pose sequence $\bm{\Psi}=[\bm{\psi}_{1},\ldots,\bm{\psi}_{T}]\in\mathbb{R}^{T\times 3}$, the VQ-VAE encoder, which consists of 1D convolutional layers, firstly encodes it into context-aware features $\hat{\bm{Z}}\in\mathbb{R}^{T^{\prime}\times d_{\bm{z}}}$, where $T^{\prime}=T/w$ denotes the frame numbers of the downsampled features. Then, we obtain the quantized sequence $\bm{Z_{q}}$ via an element-wise quantization function $Q$ that maps each element in $\bm{Z_{q}}$ to its closest codebook entry:
$\bm{Z_{q}}=Q(\hat{\bm{Z}}):=(\underset{\bm{z_{k}}\in\mathcal{Z}}{\arg\min}\|\hat{\bm{z}}_{t}-\bm{z}_{k}\|_{2})\in\mathbb{R}^{T^{\prime}\times d_{\bm{z}}}.$ (2)
@@
The quantized sequence is then reconstructed by the VQ-VAE decoder. Three 1D convolutional layers comprise the VQ-VAE decoder.
@@
Pose GPT. With the learned discrete pose space, we build a Pose GPT network Vaswani et al. (2017); Siyao et al. (2022) that maps the speech and pose style embedding into the discrete codes of the target poses. The HuBERT feature and mel spectrogram of input speech are sent into several Transformer layers (Vaswani et al., 2017) as the speech content features. HuBERT feature conveys more linguistic information while mel spectrogram contains prosody information (Chen et al., 2022). We utilize these two features for the reason that head poses are correlated to both the speech semantics and the speech prosody (e.g., stress, tone). We assign each pose sample in the training dataset with a one-hot label as the pose style embedding. It is passed through several linear layers, concatenated with the speech content features, and then fed into the Pose GPT comprising six cross-conditional Transformer layers. The predicted codes are finally sent into the VQ-VAE decoder to obtain the continuous head pose sequence. We train this Pose GPT using the teacher-forcing training scheme with the cross-entropy loss.
@@
Semantic-Aware Pose Style Matrix. Given a speech utterance of $T$ frames and the corresponding head pose sequence, HuBERT (Hsu et al., 2021) assigns each speech frame into one of 512 clusters, while the VQ-VAE encoder transforms the pose sequence into the latent sequence $\hat{\bm{Z}}$. Thus, each frame in $\hat{\bm{Z}}$ is labeled with a cluster which contains the speech semantic information. We calculate the semantic-aware pose style matrix $S\in\mathbb{R}^{512\times d_{z}}$ for each pose sequence as:
$S_{j}=\sum_{i=1}^{T^{\prime}}\hat{\bm{Z}}_{i}\cdot\delta(L_{i},j)/\sum_{i=1}^{T^{\prime}}\delta(L_{i},j),j=[1,2,\dots,512],$ (3)
@@
where $L_{i}$ denotes the cluster label for $i$ frame of $\hat{\bm{Z}}$ and $\delta$ is the Kronecker symbol that outputs 1 when the inputs are equal else 0. This equation aggregates all the poses which have the same semantic unit of HuBERT. Since the clustering step in HuBERT models the semantic information in acoustic units and the VQ-VAE encoder models the temporal pose style information, the derived $S$ represents the pose style that considers semantics.
@@
Procedure of the Pose Adapter. The pretrain, adaptation, and inference procedures of the pose adapter are listed in Alg. 2. Note that the adaptation stage involves no parameter updating.
Algorithm 2 Pretrain, Adaptation, and Inference of Pose Adapter.
@@
1:  Pretrain: Assign each pose sequence in the source training dataset with a one-hot pose style embedding. Train the VQ-VAE and Pose GPT with source training dataset. Calculate the semantic-aware pose style matrix for each sequence in the training dataset and establish the database for all pose style matrices and their corresponding pose style embedding.
@@
2:  Adaptation: Given an unseen pair of reference speech and pose sequence, calculate the semantic-aware pose style matrix for the reference data. Find the nearest pose style matrix in the established database by computing the L1 distance, and adopt the corresponding pose style embedding for inference.
@@
3:  Inference: Given a driving speech, and utilizing the retrieved pose style embedding, the pose adapter generates head poses with the pre-trained Pose GPT and the VQ-VAE decoder.